OpenStudy (anonymous):

Find k such that 9^x = e^(k x) for all values of x. I found K =LN(9) BUT,Then use your result to find the derivative of y = 9^x = e^(k x)

OpenStudy (anonymous):

i found y'= 9^x=e^((ln9)*x) but dont know if its right

OpenStudy (anonymous):

what is your y'

OpenStudy (anonymous):

9^x=e^((ln9)*x)

OpenStudy (anonymous):

thats my answer

OpenStudy (anonymous):

thats just the originalequation with k replaced so you needto fid the derivative of that

OpenStudy (anonymous):

i just plugged in ln(9) for k

OpenStudy (anonymous):

oh ok how do i find the derivative then?

OpenStudy (anonymous):

derivative of e^(some constatn*x)=some constant*e^(some constant * x)

OpenStudy (anonymous):

product rule?

OpenStudy (anonymous):

im kinda lost

OpenStudy (anonymous):

the exponent rule

OpenStudy (anonymous):

for when x is the exponent

OpenStudy (anonymous):

could u write it out?

OpenStudy (anonymous):

ln(9)*e^(x*ln(9))

OpenStudy (anonymous):

so what is e?

OpenStudy (anonymous):

the deriv of ln(9) is what?

OpenStudy (anonymous):

e is a very special non-repeating constant number

OpenStudy (anonymous):

derivative of ln(9) would be 0 since its a constant

OpenStudy (anonymous):

so it would be just x

OpenStudy (anonymous):

or e^x =1

OpenStudy (anonymous):

c=some constant number the derivative of e^(cx)=c*e^(cx)

OpenStudy (anonymous):

whats the constant number?

OpenStudy (anonymous):

ln(9) is a constant number. A constant is anything in the equation tha doesnt contain a variable

OpenStudy (anonymous):

ok